OpenStudy (anonymous):

Find the difference. 4d2e - d2e

OpenStudy (gorv):

both term has d2e

OpenStudy (gorv):

lets think d2e is not there

OpenStudy (gorv):

we left with 4-1 4-1=3 now attach d2e with it

OpenStudy (gorv):

=3d2e is our ans

OpenStudy (anonymous):

@grov that's wrong. sorry but it's just 3d

OpenStudy (anonymous):

woops @gorv that's wrong. sorry but it's just 3d

OpenStudy (gorv):

hahahahaha.....woopss that will be 3d2e

OpenStudy (gorv):

d2 means d^2

OpenStudy (gorv):

@mason437 i think u forgot 2 and e

OpenStudy (anonymous):

it (4d)(2e)-(d)(2e)

OpenStudy (anonymous):

it's *

OpenStudy (gorv):

sooooo

OpenStudy (gorv):

u cannot cancel 2e by 2e they are ni multiplication

OpenStudy (gorv):

4d means 4*d bro

OpenStudy (anonymous):

yes but 2e-2e=0

OpenStudy (gorv):

lollll..and d

OpenStudy (gorv):

so that means 43 -3 =4??

OpenStudy (gorv):

or 43-4=3?? bcozz 4-4=0

OpenStudy (anonymous):

ok then what is 4x2y-x2y?

OpenStudy (anonymous):

@gorv ok then what is 4x2y-x2y?

OpenStudy (gorv):

3x2y

OpenStudy (gorv):

ok tell me 4x^2-x^2=??

OpenStudy (anonymous):

3x^2. but were is the y?

OpenStudy (anonymous):

the d is not ^2

OpenStudy (anonymous):

how about i put this in my scientific calculator and see what i get.